OpenStudy (anonymous):

Two functions f(x) g(x) are called asymptotic f(X) g(x)-->1 as x-->(infiniti sign). Find two polynomials which are asymptotic. (Hint : think about horizontal asympototes. how can you get a horizontal aymptote of 1?)

OpenStudy (tkhunny):

Answer that last question.

OpenStudy (anonymous):

what question?

OpenStudy (tkhunny):

how can you get a horizontal aymptote of 1?)

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

?

OpenStudy (tkhunny):

Have you studied Rational Functions at all?

OpenStudy (anonymous):

a little bit , i understand a bit

OpenStudy (tkhunny):

Degree of Numerator Less Than Degree of Denominator Horizontal Asymptote, y = 0 \(y = \dfrac{1}{x+1}\) \(y = \dfrac{x-4}{x^{2}-1}\) Degree of Numerator Equal To Degree of Denominator Horizontal Asymptote, y = (Ratio of Leading Coefficients) \(y = \dfrac{2x}{x+1}\) -- y = 2/1 or just y = 2 \(y = \dfrac{x^{2}-4}{3x^{2}-1}\) -- y = 1/3 Now, I ask again, how do you get an horizontal asymptote of y = 1?

OpenStudy (anonymous):

is it like x^2 , over x^2?

OpenStudy (tkhunny):

You have it! Or (3x)/(3x) or (9x^4)/(9x^4). As long as the Degree is the same and the leading coefficient is the same. Excellent work.

OpenStudy (anonymous):

oh okay, so does that answer the question tho?

OpenStudy (tkhunny):

Of course, those are pretty boring. You may wish to dress them up a bit \[y = \dfrac{\sqrt{2}x^{2} - 5x + 11}{\sqrt{2}x^{2} + 3x - 7}\]

OpenStudy (anonymous):

where did the square root come from tho?

OpenStudy (tkhunny):

The question asks for two polynomials. I just pulled them out of a hat. ALL I care about it making the degree the same (2) and making the leading coefficient the same. \(\sqrt{2}\) may not be pretty, but as long as it's the same for both, we're good.

OpenStudy (anonymous):

oh so it can be any kind then?oh oh okay just to where we have the 1 coefficient?

OpenStudy (tkhunny):

Degree the same Leading coefficient the same Horizontal Asymptote: y = 1 Every time! Nothing else matters for the location of the horizontal asymptote.

OpenStudy (anonymous):

oh okay, so all those words are just there to confuse me then?

OpenStudy (tkhunny):

"Find two polynomials which are asymptotic" -- Pick any two you like. You don't have to use mine.

OpenStudy (anonymous):

um 4x^3+2x^2+x/ 4x^3+x^2+x ?

OpenStudy (tkhunny):

Yes, but very bad notation: (4x^3+2x^2+x)/ (4x^3+x^2+x ) The parentheses are NOT optional. Do you know why?

OpenStudy (anonymous):

yea, i coulndt think of another one . um not really

OpenStudy (tkhunny):

Order of Operations. x + 1 / x + 2 = \(x + \dfrac{1}{x} + 2\) (x + 1) / (x + 2) = \(\dfrac{x+1}{x+2}\) It is a significant difference, wouldn't you say?

OpenStudy (anonymous):

yeah, i see the difference

OpenStudy (tkhunny):

Very good. Be careful with that and move on to the next problem.

OpenStudy (anonymous):

okay thank you very much